If f(x) = 4x2 − x + 2,
find the following:

f(a + 1) =
[f(a)]2 =
f(a + h) =

Answer by Alan3354(69443)

If f(x) = 4x2 − x + 2,
find the following:
==================
f(a + 1)
f(x) = 4x^2 - x + 2
Wherever there's x, replace it with a+1
f(a+1) = 4(a+1)^2 - (a+1) + 2
= 4a^2 + 8a + 4 - a - 1 + 2
= 4a^2 + 7a + 5
=========================
Do the same way.
Use ^ (Shift 6) for exponents.
======================================
[f(a)]2 =
f(a + h) =